# New Advances in Phonons: From Band Topology to Quasiparticle Chirality

### v1

@@ VI.1.2 Topological Phonons Observed by EELS
Before 2023, experimental studies on topological phonons has remained restricted to select 3D single crystals analyzed via inelastic X-ray or neutron scattering techniques. Crucially, experimental confirmation of topological phonons in 2D materials has been lacking. Hereafter, Li et al. (2023a) employ surface sensitive high-resolution electron energy loss spectroscopy (HR-EELS) to directly map the phonon spectra of atomically thin graphene across its entire 2D Brillouin zone, as shown in Fig. 38. By comparing the experimentally determined phonon dispersion with the DFT calculations in the entire 2D Brillouin zone, Li et al. (2023a) identified two nodal-ring phonons and four Dirac phonons, as shown in Fig. 38 (c). Figure 38 (d) further show the 3D mapping of phonon spectra, which integrats 2D momentum space and energy dimensions, establishes a transformative framework for systematically identifying topological phononic states.
@@
Figure 38: Phonon spectra and experimental setup for graphene. (a) Phonon bands obtained by first-principle calculations. Topological nodal-ring phonons (NPs, green circles) and Dirac phonons (DPs, blue circles) are highlighted. (b) Schematic illustration of the high-resolution electron energy loss spectroscopy (HR-EELS) experimental setup. (c) Experimental phonon spectra of graphene along the same high-symmetry paths as (a), with NPs and DPs marked by green and blue arrows, respectively. (d) 3D HR-EELS mapping of graphene phonon spectra, integrating momentum-space and energy-resolved measurements. Adapted from Ref. Li et al. (2023a).
@@
While topological phonons in the 3D and 2D bulk state have been extensively studied, the bulk-edge correspondence of topological phonons remain to be experimentally established. For this purpose, surface and local probes of phonon excitations are required. The ability to reveal the phonon spectrum of graphene in large momentum space making HR-EELS a promising tool to uncover phononic surface state in 3D systems Zhang et al. (2018a); Miao et al. (2018); Zhang et al. (2019b). The phononic surface mode may also be encoded in electronic spectral functions through quasiparticle interactions. For instance, the drumhead surface phonon mode has been predicted in MoB2 and high-$T_{c}$ conventional superconductor MgB2. Due to the strong electron-phonon coupling, the topological phonon edge mode can couple with the conducting electrons on the surface, giving rise to electronic kinks that can be revealed using angle-resolved photoemission spectroscopy.
VI.2 Chiral Phonons
@@
As we discussed in previous sections and showed in Fig. 33, topological and chiral phonons are deeply connected in the context of Weyl phonons Zhang et al. (2025b). In this respect, the observations of Weyl phonons and chiral phononic wave functions Miao et al. (2018); Li et al. (2021a) also represent evidence of chiral phonons with circularly polarization. On the other hand, since chiral phonons are defined on AM of QPs, they are naturally connected to magnetism Kim et al. (2023); Fransson (2023); Li et al. (2024); Ren et al. (2021); Saparov et al. (2022); Peters et al. (2022); Hu et al. (2021); Xue et al. (2025); Juraschek et al. (2022); Cheng et al. (2020a); Baydin et al. (2022); Wu et al. (2023); Hernandez et al. (2023); Nova et al. (2017); Mankovsky et al. (2022); Che et al. (2024); Yang et al. (2024b), thermal Hall effect Qin et al. (2012); Kasahara et al. (2018); Grissonnanche et al. (2019); Zhang et al. (2019c); Li et al. (2020d); Grissonnanche et al. (2020); Chen et al. (2020, 2022); Li et al. (2023b); Ohe et al. (2024) and light-matter scattering processes Zhang and Niu (2015); Zhu et al. (2018); Ishito et al. (2023a, b); Zhang and Murakami (2022); Zhang et al. (2023b, 2025a, 2025b); Bourgeois et al. (2025); Ueda et al. (2023); Che et al. (2024); Yang et al. (2024b) and hence attracted significant interests in the community. In this section, we discuss recent experimental progresses that are specifically related to chiral phonons.
VI.2.1 Experimental Signature of Phonon PAM
The circular dichroism (CD) in the light-matter scatterings has been used to probe chiral phonons. Due to the PAM conservation, the microscopic light-matter scattering processes involving chiral phonons display large CD effect. Interestingly, it has been found that Zhang and Niu (2015); Zhu et al. (2018); Ishito et al. (2023a, b); Zhang and Murakami (2022); Zhang et al. (2023b, 2025a, 2025b); Bourgeois et al. (2025) the light-matter scattering can display new selection rules when the incident photon momentum, $k_{i}$, and scattered photon momenutm, $k_{f}$, are along the rotational axis. Figure 40 (a) shows Stokes and anti-Stokes Raman spectra of a truly chiral phonon system $\alpha$-HgS. The red and blue curves correspond to the $\Gamma_{3}^{(2)}$ doublet phonon modes and are obtained in ($\epsilon_{i}$, $\epsilon_{f}$)=(L, R) and (R, L) polarization conditions, respectively. Here $L$ and $R$ stand for circular left and circular right polaritons with photon PAM $\sigma_{L}$=$-1$ and $\sigma_{R}$=$+1$. Notably, the PAM conservation for a $C_{3}$-symmetric direction enforces a selection rule: $\sigma_{f}$--$\sigma_{i}$=$-l_{q}^{ph}$ $mod$ 3, in agreement with experimental observations. The chiral-phonon-driven CD has also been observed in a resonant inelastic x-ray scattering (RIXS) study of chiral quartz crystals as shown in Fig. 40 (b). It is important to note that the CD in RIXS is not a direct consequence of PAM conservation as $k_{i}$ and $k_{f}$ are not along the rotational axis. Instead, it arises from the coupling between chiral phonons and the RIXS intermediate states, which shows different scattering cross-section under $L$ and $R$ polarizations. Chiral phonons have also been observed in 2D using ultrafast pump-probe techniques Zhu et al. (2018); Britt and Siwick (2023); Zhang et al. (2025c). Figure 40 (c) shows the indirect optical transitions in a transition metal dichalcogenides WSe2. The absence of inversion symmetry in WSe2 lifts the degeneracy of chiral phonon modes at the $K$ and $K$’ valleys. The photon excited hole, $h_{A}(K)$, at the valence band of the $K$-valley are scattered, through its polarization dependent interactions with chiral phonons, to the spin-split conduction band, $h_{B}(K^{\prime})$, of the $K^{\prime}$-valley, yielding CD in the time domain [Fig. 40 (b2)] and a characteristic chiral phonon energy, 29$\pm$8 meV, at delay time $t$ = 0.8 ps [Fig. 40 (b3)].
@@ VI.2.1 Experimental Signature of Phonon PAM
Figure 39: (a) Phonon dispersions near the charge-2 Dirac point in MnSi. Solid color indicates energy integration range (30–32 meV). (b) Pseudospin texture enclosing the R point. (c) INS intensity along the red circle in (b). (d) and (e) INS intensities over $\mathbf{q}$ spheres enclosing the momentum of (2.5, 2.5, 2.5) and (2.5, 3.5, 3.5), respectively. (e) and (f) Fitted-force-constant model calculations according to Eq. (94), for the same $\mathbf{q}$ spheres as in (d) and (e), respectively. Adapted from Ref. Jin et al. (2022)
@@
Although CD provides important signature of chiral phonons, it should be emphasized that CD is sensitve to chirality rather than AM. It thus calls for experimental protocals to directly reveal the AM of phonons. Historically, the AM of microscopic quantum objects, such as the AM of electron spin, has been determined through the Einstein-de Hass (EdH) effect Einstein (1915); Gans (1916); Frenkel (1979): due to total AM conservation, the suspended ferromagnet displays a macroscopic mechanical rotation under external magnetic field to compensate the total AM of microscopic spins. Motivated by this concept, it is proposed that the AM of chiral phonon can induce mechanical rotations under a thermal gradient that breaks the time-reversal symmetry Hamada et al. (2018). In the equilibrium condition, the phonon-AM can be expressed as:
@@
$n_{\sigma}(\mathbf{q})=f_{0}(\omega_{\sigma}(\mathbf{k}))-t_{0}(v_{\sigma}(\mathbf{k})\cdot\nabla T)\frac{\partial f_{0}(\omega_{\sigma}(\mathbf{k}))}{\partial T},$ (96)
@@
Fundamentally, Eq. (95)-(97) is similar to the Edelstein effect in electronic systems Einstein (1915); Gans (1916); Frenkel (1979). The presence of thermal gradient will induce mechanical torque, $\tau\sim\frac{\delta J_{ph}}{t_{0}}$. Recently, using a cantilever-based method, the phonon AM induced mechanical torque has been observed in chiral single crystal Te Zhang et al. (2024a). The experimentally determined $\tau\sim 5\times 10^{-11}N\cdot m$ is remarkably consistent with theoretical calculations, supporting phonon AM in chiral crystals.
Figure 40: Anti-Stokes (a1) and Stokes (a2) spectra of a doubly degenerate chiral phonon modes in $\alpha$-HgS. Chiral optical phonons slightly away from the $\Gamma$ point with PAM=+1 (red) and -1 (blue) are selectively probed under ($\hat{\varepsilon}_{i}$, $\hat{\varepsilon}_{f}$)=($L,R$) and ($R,L$) polarization conditions. Here, $L$ and $R$ stand for left and right circular polarization, respectively Ishito et al. (2023a). (b1)-(b3) show chiral phonon driven circular dichroism (CD) in the indirect optical transitions in WSe2 Zhu et al. (2018). $h_{A}(K)$ and $h_{B}(K^{\prime})$ represent holes at the valence band of the $K$ valley and the spin-split band of the $K^{\prime}$ valley, respectively. The experimentally determined CD at 82 K and transient CD acquired at delay time, $\tau$=0.8 ps [(b2)], are used to estimate a chiral phonon energy 29$\pm$8 meV [(b3)], in agreement with first principles calculations. (c1)-(c2) show experimental signature of a chiral-phonon-activated spin Seebeck effect Kim et al. (2023). Applying transient thermal gradient on left- and right- handed materials [(c1)] yields spin currents with a $\pi$-phase difference [(c2)]. Adapted from Ref. Ishito et al. (2023b); Zhu et al. (2018); Kim et al. (2023)
@@
VI.2.2 Experimental Evidence of Chiral Phonons Coupled with Magnetism
@@
In real space, chiral phonons correspond to collective circular motions of charged ions. Therefore, chiral phonons can carry finite magnetic moment on the order of $\mu_{ph}\sim e\hbar/(2m_{ion})$, where $m_{ion}$ is the ionic mass that is much larger than the electron mass, $m_{e}$, yielding $\mu_{ph}/\mu_{B}\sim 10^{-3}-10^{-5}$. Surprisingly, recent experimental studies have shown giant phonon magnetic moments, $g_{ph}\mu_{ph}\sim g\mu_{B}$, in various semimetals and semiconductors, including CeCl3, Cd3As2, PbTe, and Fe2Mo3O8 Juraschek et al. (2022); Cheng et al. (2020a); Baydin et al. (2022); Wu et al. (2023). To understand these observations, Ren et al. pictured the phonon magnetic moment as electronic orbital magnetization in an adiabatic response to the underlying ionic circular motions Ren et al. (2021). They reformulated the phonon magnetism by introducing a momentum $k$-dependent Born effective charge and discovered a topological magnetization term, which arises from the chiral phonon-modified electronic energy with non-trivial Berry curvature. In topological semimetals and insulators, integration of the Berry curvature near the Yang’s monopole gives rise to giant phonon magnetic moment, $g_{ph}\mu_{ph}\sim g\mu_{B}$. Interestingly, a recent time-domain spectroscopy study of $Pb_{1-x}Sn_{x}Te$ films showed that crossing the topological insulator to trivial insulator phase boundary yields a 3-order of magnitude reduction of $g_{ph}\mu_{ph}$ in agreement with theoretical predictions.
In the presence of spin-orbital coupling, chiral phonons induce spin magnetization via dynamical modulation of electronic states. It is theoretically demonstrated Hamada and Murakami (2020a); Yao and Murakami (2024) that due to the time-dependent Hamiltonian by the chiral phonons, the expectation value of the spin against a Berry-phase term, and becomes nonzero in the presence of chiral phonons. Similarly, in magnets with localized spins, it is theoretically shown that the chiral phonons change the total magnetization by the Berry phse term in the magnon population Yao and Murakami (2024).
The coupling between chiral phonon and magnetism has also been suggested in ultrafast optical studies of nonequilibrium states. Figure 40 (c) shows an experimental observation of a chiral-phonon-activated spin Seebeck effect (CPASS). Shining an ultrafast laser pulse builds a transient temperature gradient across the heterostructure and excite chiral phonons from the non-magnetic 2D hybrid organic-inorganic perovskite (HOIP). It has been argued that the AM of chiral phonons is transferred to electrons in the Cu, resulting in spin currents without external magnetic fields. The generated spin current is then probed by the time-resolved Kerr rotation angle $\Delta\theta_{kerr}$. By switching the handedness of HOIP from left (S, red) to right (R, blue), the oscillating Kerr signal in the time domain shows a $\pi$-phase shift, consistent with the chirality-dependence of the spin current. Intriguingly, Kim et al. estimated the CPASS coefficient in their setup and obtained a value that is 1$\sim$3 orders greater than the conventional spin Seeback coefficient based on magnetic materials Kim et al. (2023). Another interesting observation showing chiral phonon mediated demagnetization has been reported recently. Recently, another ultrafast diffraction study of magnetic Ni reported evidence of nearly instantaneous but long-lasting chiral phonons Tauchert et al. (2022). Stupakiewicz et al. (2021) interpreted their observations as a rapid AM transfer process between electrons and chiral phonons that occurs before the macroscopic rotation happens. This observation provides insights into a fundamental question in ultrafast demagnetization and may help to understand the time-scale of phonon Einstein–de Haas effect.
@@
Chiral phonons have broader interests in the context of quantum information science. It has been predicted the phonon AM can contribute to the phonon Einstein-de Haas effect Hamada et al. (2018). Applying heat flow through a chiral crystal breaks $\mathcal{T}$ and causes rigid body rotations. Kong et al. (2022) showed that chiral phonons can selectively light up dark excitons from different valleys of MoSi2N4 in the photoexcitation process, yielding a chiral phonon-triggered photoluminescence. Chiral phonons were also considered for non-reciprocal thermal transport Qin et al. (2012); Kasahara et al. (2018); Grissonnanche et al. (2019); Zhang et al. (2019c); Li et al. (2020d); Grissonnanche et al. (2020); Chen et al. (2020, 2022); Li et al. (2023b); Ohe et al. (2024). More recently, chiral phonons were proposed to detect dark matter through its non-gravitational interaction with standard-model matter Romao et al. (2023). Furthermore, quasiparticle interactions between topological magnon, polaron, and phonon were predicted to give rise to chiral edge phonon modes for future heat management of microelectronic devices.

### v2

@@ VI.1.2 Topological Phonons Observed by EELS
Early experimental studies on topological phonons have been focused 3D single crystals. The experimental confirmation of topological phonons in 2D materials remain unresolved until li2023direct . Using surface sensitive HR-MEELS, li2023direct mapped the phonon spectra of atomically thin graphene across its entire 2D Brillouin zone, as shown in Fig. 36. By comparing the experimentally determined phonon dispersion with DFT calculations, li2023direct identified two nodal-ring phonons and four Dirac phonons, as shown in Fig. 36 (c). Figure 36 (d) further show the 3D mapping of phonon spectra, which integrates 2D momentum space and energy dimensions.
@@
Figure 36: Phonon spectra and experimental setup for graphene. (a) Phonon bands obtained by first-principle calculations. Topological nodal-ring phonons (NPs, green circles) and Dirac phonons (DPs, blue circles) are highlighted. (b) Schematic illustration of the high-resolution electron energy loss spectroscopy (HR-EELS) experimental setup. (c) Experimental phonon spectra of graphene along the same high-symmetry paths as (a), with NPs and DPs marked by green and blue arrows, respectively. (d) 3D HR-EELS mapping of graphene phonon spectra, integrating momentum-space and energy-resolved measurements. Adapted from Ref. li2023direct .
@@
While topological phonons in the 3D and 2D bulk state have been extensively studied, the bulk-edge correspondence of topological phonons remain to be experimentally established. For this purpose, surface and local probes of phonon excitations are required. The ability to reveal the phonon spectrum of graphene in large momentum space making HR-MEELS a promising tool to uncover phononic surface state in 3D systems zhang2018double ; miao2018observation ; zhang2019phononic . The phononic surface mode may also be encoded in electronic spectral functions through quasiparticle interactions. For instance, the drumhead surface phonon mode has been predicted in MoB2 and high-$T_{c}$ conventional superconductor MgB2. Due to the strong electron-phonon coupling, the topological phonon edge mode can couple with the conducting electrons on the surface, giving rise to electronic kinks that can be revealed using angle-resolved photoemission spectroscopy.
VI.2 Circularly Polarized Phonons
@@
As we discussed in previous sections (seeFig. 31), topological andcircularly polarized phonons are deeply connected in the context of Weyl phonons zhang2025weyl . In this respect, the observations of Weyl phonons and circularly polarized phononic wave functions miao2018observation ; li2021observation represent evidence of circularly polarized phonons. On the other hand, since circularly polarized phonons are defined on AM, they are naturally connected to magnetism kim2023chiral ; fransson2023chiral ; li2024chiral ; ren2021phonon ; saparov2022lattice ; peters2022magnetic ; hu2021phonon ; xue2025extrinsic ; juraschek2022giant ; cheng2020large ; baydin2022magnetic ; wu2023fluctuation ; hernandez2023observation ; nova2017effective ; mankovsky2022angular ; che2025magnetic ; yang2025inherent , thermal Hall effect qin2012berry ; kasahara2018majorana ; grissonnanche2019giant ; zhang2019thermal ; li2020phonon ; grissonnanche2020chiral ; chen2020enhanced ; chen2022large ; li2023phonon ; ohe2024chirality and light-matter scattering processes zhang2015chiral ; zhu2018observation ; ishito2023chiral ; ishito2023truly ; zhang2022chiral ; zhang2023weyl ; zhang2025thechirality ; zhang2025weyl ; bourgeois2025strategy ; ueda2023nature ; che2025magnetic ; yang2025inherent ; li2024phonon . For this reason, circularly polarized phonons have attracted significant interests in the community. In this section, we highlight some recent experimental progresses that are specifically related to circularly polarized phonons.
VI.2.1 Experimental Signature of Phonon PAM and AM
[…]
@@ VI.2.1 Experimental Signature of Phonon PAM and AM
Figure 37: (a) Phonon dispersions near the charge-2 Dirac point in MnSi. Solid color indicates energy integration range (30–32 meV). (b) Pseudospin texture enclosing the R point. (c) INS intensity along the red circle in (b). (d) and (e) INS intensities over $\mathbf{q}$ spheres enclosing the momentum of (2.5, 2.5, 2.5) and (2.5, 3.5, 3.5), respectively. (e) and (f) Fitted-force-constant model calculations according to Eq. (102), for the same $\mathbf{q}$ spheres as in (d) and (e), respectively. Adapted from Ref. jin2022chern
@@
Although CD provides important signature of circularly polarized phonons, it should be emphasized that CD is sensitve to mirror-symmetry breaking rather than the AM. It thus calls for experimental protocals to directly reveal the AM of phonons. Historically, the AM of microscopic quantum objects, such as the AM of electron spin, has been determined through the Einstein-de Hass (EdH) effect einstein1915experimental ; gans1916paramagnetismus ; frenkel1979history . Due to total AM conservation, the suspended ferromagnet displays a macroscopic mechanical rotation under external magnetic field to compensate the total AM of microscopic spins. Motivated by this concept, it is proposed that the AM of circularly polarized phonon can induce mechanical rotations under a thermal gradient that breaks the time-reversal symmetry hamada2018phonon . In the equilibrium condition, the phonon-AM can be expressed as:
@@
$n_{\sigma}(\mathbf{k})=f_{0}(\omega_{\sigma}(\mathbf{k}))-t_{0}(v_{\sigma}(\mathbf{k})\cdot\nabla T)\frac{\partial f_{0}(\omega_{\sigma}(\mathbf{k}))}{\partial T},$ (104)
@@
Fundamentally, Eq. (103)-(105) is similar to the Edelstein effect in electronic systems einstein1915experimental ; gans1916paramagnetismus ; frenkel1979history . The presence of thermal gradient will induce mechanical torque, $\tau\sim\frac{\delta J_{ph}}{t_{0}}$. Recently, using a cantilever-based method (Fig. 39), the phonon AM induced mechanical torque has been observed in chiral single crystal Te zhang2024observation . The experimentally determined $\tau\sim 5\times 10^{-11}N\cdot m$ is remarkably consistent with theoretical calculations, supporting phonon AM in chiral crystals.
[…]
Figure 39: (a) and (b), Photos of experimental set-ups for the left- and right-handed Te single crystals. (c) Torque responses for the left-handed Te single crystal. Turning on the heating laser drives opposite sign of torque $\tau$ from the two cantilevers over multiple cycles. (d), The same measurement for the right-handed Te single crystal. Adapted from Ref. zhang2024observation
@@
VI.2.2 Experimental Evidence of Circularly polarized Phonons Coupled with Magnetism
@@
In real space, circularly polarized phonons correspond to collective circular motions of charged ions. Therefore, circularly polarized phonons can carry finite magnetic moment on the order of $\mu_{ph}\sim e\hbar/(2m_{ion})$, where $m_{ion}$ is the ionic mass that is much larger than the electron mass, $m_{e}$, yielding $\mu_{ph}/\mu_{B}\sim 10^{-3}-10^{-5}$. Surprisingly, recent experimental studies have shown giant phonon magnetic moments, $g_{ph}\mu_{ph}\sim g\mu_{B}$, in various semimetals and semiconductors, including CeCl3, Cd3As2, PbTe, and Fe2Mo3O8, Co3Sn2S2, etc. cheng2020large ; baydin2022magnetic ; wu2023fluctuation ; luo2023large ; basini2024terahertz ; davies2024phononic ; che2025magnetic ; yang2025inherent ; lujan2024spin ; mustafa2025origin . To understand these observations, Ren et al. pictured the phonon magnetic moment as electronic orbital magnetization in an adiabatic response to the underlying ionic circular motions ren2021phonon . They reformulated the phonon magnetism by introducing a momentum $k$-dependent Born effective charge and discovered a topological magnetization term, which arises from the circularly polarized phonon-modified electronic energy with non-trivial Berry curvature. In topological semimetals and insulators, integration of the Berry curvature near the Yang’s monopole gives rise to giant phonon magnetic moment, $g_{ph}\mu_{ph}\sim g\mu_{B}$. Interestingly, a recent time-domain spectroscopy study of $Pb_{1-x}Sn_{x}Te$ films showed that crossing the topological insulator to trivial insulator phase boundary yields a 3-order of magnitude reduction of $g_{ph}\mu_{ph}$ in agreement with theoretical predictions. Moreover, two independent groups have successfully reproduced the circularly polarized phonon splitting in Co3Sn2S2 using first-principles calculations: wang2025abinitiotheorylarge introduced a Raman term arising from the differences in phonon self-energies between distinct spin channels, whereas zhang2025electronic applied the molecular Berry curvature framework. Despite different approaches, consistent results have been obtained. This consistency suggest that the observed phonon splitting originates from symmetry breaking induced by the intrinsic magnetic order. Further recent theoretical developments on phonon magnetic moments can be found in Refs. wang2025abinitiotheorylarge ; juraschek2022giant ; geilhufe2023electron ; zhang2023gate ; chaudhary2024giant ; chaudhary2025anomalous ; zhang2025electronic .
[…]
@@
Circularly polarized phonons have shown application potentials in quantum information science chen2019entanglement . A recent study from kong2022comprehensive showed that circularly polarized phonons can selectively light up dark excitons from different valleys of MoSi2N4 in the photoexcitation process, yielding a circularly polarized phonon-triggered photoluminescence. Circularly polarized phonons have also been considered for non-reciprocal thermal transport qin2012berry ; kasahara2018majorana ; grissonnanche2019giant ; zhang2019thermal ; li2020phonon ; grissonnanche2020chiral ; chen2020enhanced ; chen2022large ; li2023phonon ; ohe2024chirality . More recently, circularly polarized phonons have been proposed to detect dark matter through its non-gravitational interactions with standard-model matter romao2023chiral . Furthermore, quasiparticle interactions between topological magnon, polaron, and phonon were predicted to give rise to chiral edge phonon modes for future heat management of microelectronic devices.